# A family of log-correlated Gaussian processes

### v1

@@ 2.1. An extension to a generalized stable process
${\mathbb{E}}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{\alpha}$ $\leq\left({\mathbb{E}}\left|\int_{0}^{\infty}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(t)\right)f(t){\rm d}t\right|^{2}\right)^{\alpha/2}$
$=\left(2\Psi_{|f|}(r)\right)^{\alpha/2}\leq C\left(r\wedge r^{-1}\right)^{\alpha/2},$
and hence (2.13) follows. ∎
@@ 3. A central limit theorem
We prove Theorem 1.4. Since all $G_{n}(f)$ and $G(f)$ are linear in $f$, it suffices to show
@@
for all $f\in{\mathcal{F}}_{0}$. We start by evaluating ${\mathbb{E}}e^{{\rm i}G_{n}(f)}$. We first introduce some notations. Let $(\eta_{j}^{(r)})_{j\in\mathbb{N}}$ denote i.i.d. Bernoulli random variables with parameter $r\in(0,1)$, and set $\tau_{j}^{(r)}:=\sum_{k=1}^{j}\eta_{k}^{(r)}$. Introduce
@@
This is because for $f\in{\mathcal{F}}_{0}$ we have for all $r>0$,
@@
First, we re-state the point-process convergence in (3.7) as the convergence of a sequence of random points. Set $\widehat{j}_{n,1}:=\min\{j\in\mathbb{N}:\eta_{j}^{(r/n)}=1\}$, and then iteratively
@@
$8\sum_{j=1}^{\left\lfloor n/r\right\rfloor}\left|f_{n,j}\right|\sum_{k=0}^{\infty}\left|f_{n,j+k}\right|\frac{1}{4}\left(1-\left(1-\frac{2r}{n}\right)^{2j}\right)\left(1-\frac{2r}{n}\right)^{j^{\prime}-j}$
@@
$8\sum_{j=\left\lfloor n/r\right\rfloor+1}^{\infty}\left|f_{n,j}\right|\sum_{k=0}^{\infty}\left|f_{n,j+k}\right|\frac{1}{4}\left(1-\left(1-\frac{2r}{n}\right)^{2j}\right)\left(1-\frac{2r}{n}\right)^{j^{\prime}-j}$
@@
4. Log-correlated Gaussian processes indexed by metric spaces
We introduce a family of log-correlated generalized Gaussian processes indexed by metric spaces $({\mathsf{M}},{\mathsf{d}})$ under the Assumption (1.1) and their stable extensions in a unified framework. We also provide a central limit theorem.
@@ 4.1. A stochastic-integral representation
Let $(E,{\mathcal{E}},\mu)$ and $\{A_{x}\}_{x\in{\mathsf{M}}}\subset{\mathcal{E}}$ be as in Assumption 1.1. Let ${\mathbb{M}}_{p}(E)$ denote the space of Radon point measures on $E$, and $\mathcal{M}_{p}(E)$ its Borel $\sigma$-algebra (see e.g. [39]). For each $r>0$, let $\mathbb{P}_{r}^{\prime}\equiv\mathbb{P}_{r,\mu}^{\prime}$ denote the probability measure on $({\mathbb{M}}_{p}(E),\mathcal{M}_{p}(E))$ induced by the Poisson point process on $(E,{\mathcal{E}})$ with intensity measure $r\cdot\mu$. That is, letting $m$ denote an element in ${\mathbb{M}}_{p}(E)$,
$\mathbb{P}^{\prime}_{r}(m(A)=k)\equiv\mathbb{P}^{\prime}_{r}\left(\left\{m\in\mathbb{M}_{p}(E):m(A)=k\right\}\right)=\frac{(r\mu(A))^{k}}{k!}e^{-r\mu(A)},k\in\mathbb{N}_{0},$
@@
$\widetilde{p}_{r}(x):=\mathbb{P}^{\prime}_{r}\left(m(A){\rm~{}is~{}odd}\right)=\frac{1}{2}\left(1-e^{-2r\mu(A_{x})}\right),x\in{\mathsf{M}}.$
@@
$G_{\alpha}(f):=\int_{\mathbb{R}_{+}\times{\mathbb{M}}_{p}(E)}\int_{x\in{\mathsf{M}}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(x)\right)f(x)\lambda({\rm d}x)M_{\alpha}({\rm d}r{\rm d}m),$
where $M_{\alpha}$ is a symmetric $\alpha$-stable random measure on +×$\mathbb{R}_{+}\times{\mathbb{M}}_{p}(E)$ with control measure $2r^{-1}{\rm d}r{\rm d}\mathbb{P}_{r}^{\prime})$, and $f\in{\mathcal{F}}_{0}({\mathsf{M}})$ from following assumption.
@@
We let ${\mathcal{F}}_{0}({\mathsf{M}})$ denote the class of all bounded and measurable functions $f:{\mathsf{M}}\to\mathbb{R}$ (measurable with respect to the Borel $\sigma$-algebra induced by ${\mathsf{d}}$) satisfying the following:
@@
    ${\left\|f\right\|}_{1,\delta}:=\int_{\mathsf{M}}\mu(A_{x})^{\delta}|f(x)|\lambda({\rm d}x)<\infty$ for some $\delta>0$,
@@
    $\int_{\mathsf{M}}e^{-2r\mu(A_{x})}|f(x)|\lambda({\rm d}x)\leq Cr^{-\delta^{\prime}}$ for all $r>0$ and some fixed constants $C,\delta^{\prime}>0$.
It is useful to recall that when $\mu(A_{o})=0$ for some $o\in{\mathsf{M}}$ (usually playing the role of the origin), we have $\mu(A_{x})={\mathsf{d}}(o,x)$. In particular, the above assumption can be viewed as a generalization of the assumption on $f$ in (2.3). Indeed, with ${\mathsf{M}}=\mathbb{R}_{+}=[0,\infty)$, ${\mathsf{d}}$ being the Euclidean metric, and $\lambda$ being the Lebesgue measure, the integrability condition in ${\mathcal{F}}_{0}$ in (2.3) is exactly condition (ii) above, which also implies (i) (since $f$ is bounded). Condition (iii) follows with $\delta^{\prime}=1$ since $\mu(A_{x})=x$ and $f(x)$ is bounded.
The following is the main result for log-correlated Gaussian processes and their stable extensions indexed by metric spaces, where $\Gamma(x,y)$ is the same as in Theorem 1.2.
@@
For each $\alpha\in(0,2]$, $\{G_{\alpha}(f)\}_{f\in{\mathcal{F}}_{0}({\mathsf{M}})}$ is a family of symmetric $\alpha$-stable random variables with
(4.1) ${\mathbb{E}}\exp\left({\rm i}G_{\alpha}(f)\right)=\exp\left(-2\int_{0}^{\infty}{\mathbb{E}}^{\prime}_{r}\left|\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(x)\right)f(x)\lambda({\rm d}x)\right|^{\alpha}\frac{{\rm d}r}{r}\right).$
@@
(4.2) ${\rm{Cov}}(G_{2}(f),G_{2}(g))=\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)g(y)\Gamma(x,y)\lambda({\rm d}x)\lambda({\rm d}y),$
for all $f,g\in{\mathcal{F}}_{0}({\mathsf{M}})$, where
(4.3) $\Gamma(x,y)=\log\frac{\mu(A_{x})+\mu(A_{y})}{{\mathsf{d}}(x,y)}=4\int_{0}^{\infty}\frac{\widetilde{\Gamma}_{r}(x,y)}{r}{\rm d}r,$
@@
(4.4) $\widetilde{\Gamma}_{r}(x,y):={\mathbb{E}}_{r}^{\prime}\left(\left({{\bf 1}}_{\left\{m(A_{x}){\rm~{}is~{}odd}\right\}}-\mathbb{P}_{r}^{\prime}(m(A_{x}){\rm~{}is~{}odd})\right)\left({{\bf 1}}_{\left\{m(A_{y}){\rm~{}is~{}odd}\right\}}-\mathbb{P}_{r}^{\prime}(m(A_{y}){\rm~{}is~{}odd}\right)\right).$
@@
(4.5) $\Psi_{\alpha,f}(r)={\mathbb{E}}^{\prime}_{r}\left|\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~{}is~{}odd}\right\}}-\widetilde{p}_{r}(A_{x})\right)f(x)\lambda({\rm d}x)\right|^{\alpha}.$
@@
(4.6) $\int_{0}^{\infty}\frac{\Psi_{\alpha,f}(r)}{r}{\rm d}r<\infty.$

### v3

@@ 2.2. A representation for the bi-fractional Brownian motion with $H\in(0,1/2)$
$\widetilde{G}^{H,K}(t)=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}t){\rm~is~odd}\right\}}-\widetilde{p}_{2H,r}(t)\right)\widetilde{M}_{2,K}({\rm d}r{\rm d}\omega^{\prime}),$
where this time $\widetilde{M}_{2,K}$ is on $\mathbb{R}_{+}\times\Omega^{\prime}$ with control measure $C_{K}r^{-K-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$ where $\mathbb{P}^{\prime}$ is a probability measure on $\Omega^{\prime}$, under which $N^{\prime}$ is a standard Poisson process and $\mathcal{S}_{2H,r}$ is such that ${\mathbb{E}}^{\prime}e^{-\theta\mathcal{S}^{\prime}_{2H,r}}=e^{r\theta^{2H}}$ for all $\theta\geq 0$, $N^{\prime}$ and $\mathcal{S}_{2H,r}^{\prime}$ are independent, and
$\widetilde{p}_{2H,r}(t):=\mathbb{P}^{\prime}\left(N^{\prime}(\mathcal{S}_{2H,r}^{\prime}t){\rm~is~odd}\right)={\mathbb{E}}^{\prime}\left(\frac{1}{2}\left(1-e^{-2\mathcal{S}^{\prime}_{2H,r}t}\right)\right)=\frac{1}{2}\left(1-e^{-(2t)^{2H}r}\right).$
Lemma 2.4.
Under the notations above, for all $H\in(0,1/2]$ and $K\in(0,1)$ we have
${\rm{Cov}}\left(\widetilde{G}^{H,K}(s),\widetilde{G}^{H,K}(t)\right)=\frac{2^{(2H-1)K}}{4}\left((s^{2H}+t^{2H})^{K}-(t-s)^{2HK}\right),0\leq s<t<\infty.$
Proof.
Assume $s<t$. We first compute (with respect to $\mathbb{P}^{\prime}$)
${\rm{Cov}}^{\prime}\left({{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}s){\rm~is~odd}\right\}},{{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}t){\rm~is~odd}\right\}}\right)=\frac{1}{4}\left(e^{-2^{2H}(t-s)^{2H}r}-e^{-2^{2H}(s^{2H}+t^{2H})r}\right).$
Then,
${\rm{Cov}}\left(\widetilde{G}^{H,K}(s),\widetilde{G}^{H,K}(t)\right)$ $=C_{K}\int_{0}^{\infty}r^{-K-1}{\rm d}r{\rm{Cov}}^{\prime}\left({{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}s){\rm~is~odd}\right\}},{{\bf 1}}_{\left\{N^{\prime}(\mathcal{S}_{2H,r}^{\prime}t){\rm~is~odd}\right\}}\right)$
$=\frac{C_{K}}{4}\int_{0}^{\infty}\left(e^{-2^{2H}(t-s)^{2H}r}-e^{-2^{2H}(s^{2H}+t^{2H})r}\right)r^{-K-1}{\rm d}r$
$=\frac{C_{K}}{4}\frac{\Gamma(1-K)}{K}2^{2HK}\left(\left(s^{2H}+t^{2H}\right)^{K}-(t-s)^{2HK}\right).$
Note that when $H=1/2$, this representation was known [9, 14]. A new idea here for $H\in(0,1/2)$ is the introduction of a stable subordinator $\mathcal{S}_{2H,r}^{\prime}$ in the representation, and the same idea shall be applied in Section 4.
Remark 2.5.
In fact, one can have accordingly stochastic representations [14] for bi-fractional Brownian motions indexed by metric spaces. Furthermore, replacing Gaussian random measures by symmetric $\alpha$-stable ones, we obtain immediately extensions of bi-fractional Brownian motions, which we may name bi-fractional stable motions. We omit the details.
Remark 2.6.
Our result can be interpreted as a limit theorem for ${\mathbb{B}}^{H}$ as $H\downarrow 0$ with a different shift from the one considered by Neuman and Rosenbaum, [37] in (1.6). Our choice of the shift is in fact related to the decomposition of the fractional Brownian motion by a bi-fractional Brownian motion due to Lei and Nualart, [30]. Namely, ${\mathbb{B}}^{H}$ satisfies
(2.15) $\left({\mathbb{B}}^{H}_{t}\right)_{t\in[0,\infty)}\stackrel{{\scriptstyle d}}{{=}}\left(G^{1/2,2H}_{t}+W^{H}_{t}\right)_{t\in[0,\infty)},$
where $G^{1/2,2H}$ is a bi-fractional Brownian motion introduced earlier and $W^{H}$ satisfies
${\rm{Cov}}\left(W^{H}_{s},W^{H}_{t}\right)=\frac{1}{2}\left(s^{2H}+t^{2H}-(s+t)^{2H}\right),$
and the two processes are independent. Moreover, the decomposition (2.15) can hold in the almost sure sense. Let $(\Omega,{\mathcal{F}},\mathbb{P})$ be the by-default probability space that the random measures are defined on. Set $E=\mathbb{R}_{+}\times\Omega^{\prime}$ where $(\Omega^{\prime},{\mathcal{F}}^{\prime},\mathbb{P}^{\prime})$ is another probability space on which a standard Poisson process $N^{\prime}$ is defined, and $\widetilde{M}_{2,H}$ as a Gaussian random measure (defined on $(\Omega,{\mathcal{F}},\mathbb{P})$) on $E=\mathbb{R}_{+}\times\Omega^{\prime}$ with control measure $C_{2H}r^{-2H-1}{\rm d}r{\rm d}\mathbb{P}^{\prime}$ with $C_{\beta}=\beta 2^{-\beta}/\Gamma(1-\beta),\beta\in(0,1)$. Then, (2.15) holds with the following:
(2.16) ${\mathbb{B}}_{t}^{H}$ $:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}{{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}\widetilde{M}_{2,2H}({\rm d}r{\rm d}\omega^{\prime}),$
(2.17) $G^{1/2,2H}_{t}$ $:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\left({{\bf 1}}_{\left\{N^{\prime}(rt){\rm~is~odd}\right\}}-\widetilde{p}_{r}(t)\right)\widetilde{M}_{2,2H}({\rm d}r{\rm d}\omega^{\prime}),$
$W^{H}_{t}$ $:=\int_{\mathbb{R}_{+}\times\Omega^{\prime}}\widetilde{p}_{r}(t)\widetilde{M}_{2,2H}({\rm d}r{\rm d}\omega^{\prime}),$
with $\widetilde{p}_{r}(t)=\mathbb{P}^{\prime}(N^{\prime}(rt){\rm~is~odd})=\left(1-e^{-2rt}\right)/2$ [9, 14].
Now, one can write
$G^{1/2,2H}_{t}={\mathbb{B}}^{H}_{t}-W_{t}^{H}.$
That is, $G^{1/2,2H}$ is ${\mathbb{B}}^{H}$ with a random shift $W^{H}$. The convergence of $H^{-1/2}G^{1/2,2H}\Rightarrow G^{1/2}$ is now a special case of (1.2).
@@ 3. A central limit theorem
Another feature of the introduced process $G\equiv G^{1/2}$ is that it arises from a simple aggregated model. The aggregated model here is a variation of the one proposed by Enriquez, [11] when investigating approximations of a fractional Brownian motion. Connections between aggregated models and fractional Gaussian processes and random fields with long-range dependence [39, 44] are well-known [4, 5, 27, 35] (in particular, [4, 5] established limit theorems for generalized processes).
We follow the notations in Shen et al., [47], where the Enriquez model was revisited and extended. The model is an aggregated model of an increasing number, denoted by $m_{n}$, of i.i.d. layers indexed by $i\in\mathbb{N}$, and each layer consists of a sequence of stationary random variables denoted by $(\mathcal{X}_{i,j})_{j\in\mathbb{N}}$ that we now introduce. First, let $\mathcal{X}$ be a standard Gaussian random variable, $q$ a uniform random variable taking values from $(0,1)$, and $\{\eta_{j}^{(q)}\}_{j\in\mathbb{N}}$ conditionally i.i.d. Bernoulli random variables with parameter $q$. Assume that $\mathcal{X}$ is independent from $q$ and $\{\eta_{j}^{(q)}\}_{j\in\mathbb{N}}$. Let $(\mathcal{X}_{i},q_{i},\{\eta_{i,j}^{(q_{i})}\}_{j\in\mathbb{N}})_{i\in\mathbb{N}}$ denote i.i.d. copies of $(\mathcal{X},q,\{\eta_{j}^{(q)}\}_{j\in\mathbb{N}})$, and set
$\tau_{i,j}^{(q_{i})}:=\sum_{k=1}^{j}\eta_{i,k}^{(q_{i})},j\in\mathbb{N}.$
Then, the $j$-th element of $i$-th layer is defined as
In order to see that the proposed model exhibits log-correlated dependence structure, a quick calculation yields
(3.1) $\lim_{n\to\infty}{\rm{Cov}}\left(\widehat{S}_{n}(s),\widehat{S}_{n}(t)\right)=\frac{1}{4}\log{\frac{s+t}{s-t}},0<s<t.$
Indeed,
$\to\frac{1}{4}\int_{0}^{\infty}\frac{1}{r}\left(e^{-2(t-s)r}-e^{-2(t+s)r}\right){\rm d}r=\frac{1}{4}\log{\frac{t+s}{t-s}}.$
The convergence in (3.1) suggests that there is no convergence of point-wise defined stochastic processes, and instead one should search for convergence in the sense of generalized processes. For this purpose, recall ${\mathcal{F}}_{\delta}$ in (2.12) and introduce
(3.2) $G_{n}(f):=2\sum_{j=1}^{\infty}\widehat{S}_{n,j}\int_{(j-1)/n}^{j/n}f(x){\rm d}x,\quad f\in{\mathcal{F}}_{\delta},$
with $\delta>0$. (Note that we do not assume $f$ to be continuous for $f\in{\mathcal{F}}_{\delta}$. If continuity is assumed, then one can replace $\int_{(j-1)/n}^{j/n}f(x){\rm d}x$ by $f(j/n)/n$ and establish the same limit theorem below.)
The second main result is the following limit theorem.
Theorem 3.1.
Assume $\lim_{n\to\infty}m_{n}/n=\infty$ and $\delta>0$. Then, for all $f\in{\mathcal{F}}_{\delta}$,
In particular, $\{G_{n}(f)\}_{f\in{\mathcal{F}}_{\delta}}\xrightarrow{\textit{f.d.d.}}\{G(f)\}_{f\in{\mathcal{F}}_{\delta}}$ as $n\to\infty$.
Remark 3.2.
We impose the condition in ${\mathcal{F}}_{\delta}$ as it is convenient for establishing our limit theorems. In particular, ${\mathcal{F}}_{\delta}$ includes all smooth functions that decay rapidly at infinity. Lévy’s continuity theorem yields immediately the convergence of $G_{n}$ in the space of tempered distributions. For the generalized process $G$ indexed by $\mathsf{M}$}, Lévy’s continuity theorem applies as long as the space of functions on $\mathsf{M}$} is a nuclear space [12, 34]. We do not pursue to strengthen the convergence in the space of distributions here.
Remark 3.3.
The process $\widehat{S}_{n,j}$ can be thought of the $H=0$ counterpart of the model considered in Enriquez, [11] and Shen et al., [47]. Therein, the density of $q$ is asymptotically equivalent to $Cr^{-2H}$ as $q\downarrow 0$ for some constant $C>0$, and the normalization $1/\sqrt{m_{n}}$ in $\widehat{S}_{n,j}$ is replaced by $1/(n^{H}\sqrt{m_{n}})$, and it was shown that $\{\widehat{S}_{n,\left\lfloor nt\right\rfloor}\}_{t\in[0,1]}\xrightarrow{\textit{f.d.d.}}c\{G^{H}_{t}\}_{t\in[0,1]}$ for some explicit constant $c$. (Therein, there is also no centering in $\widehat{S}_{n,j}$ and so the limit is actually ${\mathbb{B}}^{H}$ instead of $G^{H}$.)
3.1. Proof of Theorem 3.1
Since all $G_{n}(f)$ and $G(f)$ are linear in $f$, it suffices to show
@@
for all $f\in{\mathcal{F}}_{\delta},\delta>0$. We start by evaluating ${\mathbb{E}}e^{{\rm i}G_{n}(f)}$. We first introduce some notations. Let $(\eta_{j}^{(r)})_{j\in\mathbb{N}}$ denote i.i.d. Bernoulli random variables with parameter $r\in(0,1)$, and set $\tau_{j}^{(r)}:=\sum_{k=1}^{j}\eta_{k}^{(r)}$. Introduce
@@
This is because for $f\in{\mathcal{F}}_{\delta}$ we have for all $r>0$,
@@
First, we re-state the point-process convergence in (3.9) as the convergence of a sequence of random points. Set $\widehat{j}_{n,1}:=\min\{j\in\mathbb{N}:\eta_{j}^{(r/n)}=1\}$,
@@
$8\sum_{j=1}^{\left\lfloor n/r\right\rfloor}\left|f_{n,j}\right|\sum_{k=0}^{\infty}\left|f_{n,j+k}\right|\frac{1}{4}\left(1-\left(1-\frac{2r}{n}\right)^{2j}\right)\left(1-\frac{2r}{n}\right)^{k}$
@@
$8\sum_{j=\left\lfloor n/r\right\rfloor+1}^{\infty}\left|f_{n,j}\right|\sum_{k=0}^{\infty}\left|f_{n,j+k}\right|\frac{1}{4}\left(1-\left(1-\frac{2r}{n}\right)^{2j}\right)\left(1-\frac{2r}{n}\right)^{k}$
@@
4. Log-correlated bi-fractional Brownian motions indexed by metric spaces
We introduce a family of log-correlated generalized Gaussian processes indexed by metric spaces $({\mathsf{M}},{\mathsf{d}})$ under the Assumption 1.2 and their stable extensions in a unified framework. We also provide a central limit theorem.
@@ 4.1. A stochastic-integral representation
Let $(E,{\mathcal{E}},\mu)$ and $\{A_{x}\}_{x\in{\mathsf{M}}}\subset{\mathcal{E}}$ be as in Assumption 1.2. Let ${\mathbb{M}}_{p}(E)$ denote the space of Radon point measures on $E$, and $\mathcal{M}_{p}(E)$ its Borel $\sigma$-algebra (see e.g. [40]). For notational convenience, we write
$\beta=2H\in(0,1]$
in this section. For each $r>0$, let $\mathbb{P}_{\beta,r}^{\prime}\equiv\mathbb{P}_{\beta,r,\mu}^{\prime}$ denote the probability measure on $({\mathbb{M}}_{p}(E),\mathcal{M}_{p}(E))$ induced by the Poisson point process on $(E,{\mathcal{E}})$ with random intensity measure $\mathcal{S}_{\beta,r}\cdot\mu$, where $\mathcal{S}_{\beta,r}$ is a totally skewed $\beta$-stable random variable with
${\mathbb{E}}\exp(-\theta\mathcal{S}_{\beta,r})=\exp\left(-r\theta^{\beta}\right),\theta>0.$
(That is, $\mathcal{S}_{\beta,r}$ is a $\beta$-stable subordinator evaluated at time $r$.) Note that the special case $\mathcal{S}_{1,r}=r$ (and this section might be easier to read first assuming $\beta=1$). In particular, letting $m$ denote an element in ${\mathbb{M}}_{p}(E)$,
$\mathbb{P}^{\prime}_{\beta,r}(m(A)=k)$ $\equiv\mathbb{P}^{\prime}_{\beta,r}\left(\left\{m\in\mathbb{M}_{p}(E):m(A)=k\right\}\right)$
$={\mathbb{E}}\left(\frac{(\mathcal{S}_{\beta,r}\mu(A))^{k}}{k!}e^{-\mathcal{S}_{\beta,r}\mu(A)}\right),k\in\mathbb{N}_{0},$
@@
$\widetilde{p}_{\beta,r}(x):=\mathbb{P}^{\prime}_{\beta,r}\left(m(A){\rm~is~odd}\right)={\mathbb{E}}\left(\frac{1}{2}\left(1-e^{-2\mathcal{S}_{\beta,r}\mu(A_{x})}\right)\right)=\frac{1}{2}\left(1-e^{-(2\mu(A_{x}))^{\beta}r}\right),x\in{\mathsf{M}}.$
@@
$G_{\alpha,\beta}(f):=\int_{\mathbb{R}_{+}\times{\mathbb{M}}_{p}(E)}\int_{x\in{\mathsf{M}}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{\beta,r}(x)\right)f(x)\lambda({\rm d}x)M_{\alpha,\beta}({\rm d}r{\rm d}m),$
where $M_{\alpha,\beta}$ is a symmetric $\alpha$-stable random measure on $\mathbb{R}_{+}\times{\mathbb{M}}_{p}(E)$ with control measure $2r^{-1}{\rm d}r{\rm d}\mathbb{P}_{\beta,r}^{\prime})$, and $f\in{\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})$ defined as follows.
@@
With $\delta,\delta^{\prime}>0,\beta\in(0,1]$, we let ${\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})$ denote the class of all bounded and measurable functions $f:{\mathsf{M}}\to\mathbb{R}$ (measurable with respect to the Borel $\sigma$-algebra induced by $\mathsf{d}$}) satisfying the following:
@@
    ${\left\|f\right\|}_{1,\delta}:=\int_{\mathsf{M}}\mu(A_{x})^{\delta}|f(x)|\lambda({\rm d}x)<\infty$,
@@
    $\int_{\mathsf{M}}\int_{\mathsf{M}}e^{-{\mathsf{d}}^{\beta}(x,y)r}|f(x)f(y)|\lambda({\rm d}x)\lambda({\rm d}y)\leq Cr^{-\delta^{\prime}}$ for all $r>0$.
It is useful to recall that when $\mu(A_{o})=0$ for some $o\in{\mathsf{M}}$ (usually playing the role of the origin), we have $\mu(A_{x})={\mathsf{d}}(o,x)$. In particular, the above assumption can be viewed as a generalization of the assumption on $f$ in (2.12).
The following is the main result for log-correlated bi-fractional Brownian motions and their stable extensions indexed by metric spaces, where $\Gamma^{(\beta)}(x,y)$ is the same as in Theorem 1.3.
@@
For each $\alpha\in(0,2],\delta,\delta^{\prime}>0$, $\{G_{\alpha,\beta}(f)\}_{f\in{\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})}$ is a family of symmetric $\alpha$-stable random variables with
(4.1) ${\mathbb{E}}\exp\left({\rm i}G_{\alpha,\beta}(f)\right)=\exp\left(-2\int_{0}^{\infty}{\mathbb{E}}^{\prime}_{\beta,r}\left|\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{\beta,r}(x)\right)f(x)\lambda({\rm d}x)\right|^{\alpha}\frac{{\rm d}r}{r}\right).$
@@
(4.2) ${\rm{Cov}}(G_{2,\beta}(f),G_{2,\beta}(g))=\int_{\mathsf{M}}\int_{\mathsf{M}}f(x)g(y)\Gamma^{(\beta)}(x,y)\lambda({\rm d}x)\lambda({\rm d}y),$
for all $f,g\in{\mathcal{F}}_{\delta,\delta^{\prime},\beta}({\mathsf{M}})$, where
(4.3) $\Gamma^{(\beta)}(x,y)=\log\frac{\mu^{\beta}(A_{x})+\mu^{\beta}(A_{y})}{{\mathsf{d}}^{\beta}(x,y)}=4\int_{0}^{\infty}\frac{\Gamma_{r}^{(\beta)}(x,y)}{r}{\rm d}r,$
@@
(4.4) $\Gamma_{r}^{(\beta)}(x,y)\\ :={\mathbb{E}}^{\prime}_{\beta,r}\left(\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\mathbb{P}_{\beta,r}^{\prime}\left(m(A_{x}){\rm~is~odd})\right)\right)\left({{\bf 1}}_{\left\{m(A_{y}){\rm~is~odd}\right\}}-\mathbb{P}_{\beta,r}^{\prime}\left(m(A_{y}){\rm~is~odd}\right)\right)\right).$
@@
(4.5) $\Psi_{\alpha,\beta,f}(r)={\mathbb{E}}^{\prime}_{\beta,r}\left|\int_{\mathsf{M}}\left({{\bf 1}}_{\left\{m(A_{x}){\rm~is~odd}\right\}}-\widetilde{p}_{r}(x)\right)f(x)\lambda({\rm d}x)\right|^{\alpha}.$
@@
(4.6) $\int_{0}^{\infty}\frac{\Psi_{\alpha,\beta,f}(r)}{r}{\rm d}r<\infty.$